7. Find the general solution of the ODE below as a power series solution about the point x = 0. (15 pts. total) y+y = 0.
Answer #1

Let y = Cnx, where cn is to be determined. n=0 By taking derivatives, v = m*** ** = Žaio Ën(n − 1)(x2-2 + Žcuz =0 We can rLet us look at odd coefficients. ģi un p 5.0 - 5.43 07 = 7.5 = 7.6 57 = - -1 C C2n+1 = ( − 1) (2n + 1)! Hence, the solution